### v3

# GDPR Elements, and Future Directions

Abstract
Biometric recognition as an efficient and hard-to-forge way of identification and verification has become an indispensable part of the current digital world. The fast evolution of this technology has been a strong incentive for integration into many applications. Meanwhile, blockchain, the decentralized ledger technology, has been widely received by both research and industry in the past few years, and it is being increasingly deployed today in many different applications, such as money transfer, IoT, healthcare, or logistics. Recently, researchers have started to speculate on the pros and cons and what the best applications would be when these two technologies cross paths. This paper provides a survey of the research literature on the combination of blockchain and biometrics and includes a first legal analysis of this integration based on GDPR to shed light on challenges and potentials. Although the integration of blockchain technology into the biometric sector is still in its infancy, with a growing body of literature discussing specific applications and advanced technological setups, this paper aims to provide a holistic understanding of blockchain applicability in biometrics. Based on published studies, this article discusses, among others, practical examples combining blockchain and biometrics for novel applications in PKI systems, distributed trusted services, and identity management. Challenges and limitations when combining blockchain and biometrics that motivate future work will also be discussed; e.g., blockchain networks at their current stage may not be efficient or economical for some real-time biometric applications. Finally, we also discuss key legal aspects of the EU General Data Protection Regulation (GDPR) related to this combination of technologies (blockchain and biometrics); for example, accountability, immutability, anonymity, and data protection elements.
Index Terms:
Biometrics, Blockchain, Security, Privacy, GDPR, General Data Protection Regulation.
I Introduction
Unlike conventional authentication methods based on knowledge (e.g. password) or possession (e.g. smart card), biometric recognition relies on the concept of inherence, i.e. who someone is, which makes it more robust against fraudulent activities like forging, spoofing, etc., compared to the aforementioned methods [1]. The advantages of biometric recognition include a high level of security [2], a better user experience [3], and a fast recognition process that have allowed the use of biometric systems to improve many different applications, such as authentication systems [4], border control [5], etc.[6]. One of the technologies with great potential to be used in conjunction with biometrics is blockchain.
The advent of Distributed Ledger Technology (DLT), and in particular blockchain, has become one of the most prominent phenomena of our time. As a result, it has come to the forefront of research in many disciplines (e.g., cryptocurrency, smart contracts, identity management) aimed at including distributed data storage. Due to the success of Bitcoin and cryptocurrencies, the corresponding blockchain technology is now successfully practiced to distribute data in a secure-by-design database without the need for a central authority. Not only have the unprecedented security properties of blockchain revolutionized a wide range of financial services and digital payments, but its unique characteristics such as decentralization, immutability, auditability, fault tolerance, and availability have played a pivotal role in receiving further public attention so far.
On the one hand, it is not surprising that, due to its availability and decentralization properties, the blockchain solves the security problems arising from storing biometric features (e.g., storing biometric features of the same individual in different places belonging to independent applications). In particular, in terms of biometric verification, the so-called on-chain storage of biometric templates helps to prevent Denial of Service (DoS) attacks caused by a single point of failure. However, a major breakthrough for the combination of blockchain and biometrics would be the introduction of the Public Key Infrastructure (PKI) mechanism[7]. Owing to their inherent connection to an individual’s identity, biometrics lack PKI structure. To address this limitation, the advantages of using blockchain and more specifically a novel introduced concept called distributed identifiers (DIDs) are being investigated. Although current studies are almost focused only on the role of the blockchain in biometric verification, identity management, and e-voting, we discuss the unprecedented merits of combining blockchain and biometrics for other potential applications such as using blockchain as computing power for biometrics or introducing a foundation for PKI mechanism. In addition, we discuss the role of blockchain in providing an infrastructure for multimodal biometrics or using it as a computational mean in Proof of Work (PoW)-based consensus for biometric verification.
However, the shortcomings of combining blockchain and biometrics need to be analyzed specifically with regard to issues such as the type and location of information that can be stored on blockchain. One of the main objectives of this work is to analyze the performance of the system, for example, latency, computation cost, interoperability, and user convenience.
Another objective of this study is to provide a first legal and data protection analysis according to the General Data Protection Regulations (GDPR). Since May 25, 2018, the GDPR has been applicable in all EU member states to harmonize data privacy laws across Europe. According to the GDPR, any information related to an identified and identifiable person is defined as personal data and shall be subject to the provisions of the GDPR. Biometric data stored on the blockchain in many applications are considered personal data, which is subject to GDPR. This paper discusses the implications based on GDPR of storing biometric data on public blockchains.
In particular, the main contributions of this paper are as follows.
  • An overview of the meeting points of blockchain and biometrics with respect to both technical and legal potentials and limitations.
  • An application-based summary and categorization of the literature from the perspective of using the blockchain in favor of biometrics.
  • An in-depth analysis of typical application scenarios for combining blockchain and biometrics while taking into account well-known aspects of blockchain such as scalability, smart contracts, consensus algorithms, security and privacy [8], and legal issues.
  • A description of the unprecedented use cases for combining blockchain and biometrics and the corresponding merits.
In this regard, Section II gives a concise introduction to blockchain technology, biometrics, and their combination. In Section III, we analyze the impact of blockchain features when combined with biometrics from a technical point of view. This section discusses the prospects and limits that lie ahead of this combination. In this section, we also provide a categorization of the technical literature on combining blockchain and biometrics to improve biometric systems. Section IV contains the legal analysis based on GDPR. Section V offers some recommendations and suggestions that should be kept in mind for future studies in this line of research. Finally, Section VI concludes this survey.
II Blockchain and Biometrics: Introduction
II-A Blockchain Technology
Blockchain is a distributed ledger shared among a group of entities that are supposed to reach a consensus to register or modify transactions [9, 10, 11]. These transactions are secure, immutable, and easy to track and measure. In particular, blockchain allows entities to reach an agreement on a certain subject and records it as a transaction without the need for a central authority. The subject of transactions varies depending on the application in which the blockchain is used. The structure of a common blockchain network is shown in Fig. 1.
Figure 1: An example of a blockchain structure consisting of a continuous sequence of blocks taken from [11].
In particular, the header contains the following: the owner identifier and the identifier of the previous block, a timestamp denoting the generation time of the current block, a nonce, and a merkle tree root hash which denotes the hash value of all transactions stored in the block. The body consists of transactions and a transaction counter. The maximum number of transactions that fit within the body depends on the block size and the size of the transactions within. A digital signature based on asymmetric cryptography is used to validate transactions [11].
The most prominent characteristics of blockchain that have made it highly popular for integration with other technologies, including biometrics, are Decentralization, Immutability, Auditability, Fault tolerant, and Availability [12, 13].
Figure 2: The lifecycle of a biometric transaction on the blockchain (PT = Protected Template).
II-B Biometric Technology
Biometrics in modern computer science refers to the automated utilization of biological traits for personal identification. These traits enable humans to distinguish individuals based on their physical (e.g. hand geometry, fingerprints, face, iris, and so on) and behavioral (e.g. signature, keystroke, and etc.) characteristics. When properly implemented, they also allow computer systems to recognize patterns for security purposes [14]. The key phases of a biometric system are enrollment and authentication. These phases involve several essential processes: (i) sensor modules for modality acquisition, (ii) feature extraction in the feature module, (iii) comparison of extracted features with the database in the matching module, and (iv) template verification in the decision module. At the sensor level, biometric modalities are captured as images, videos, audio, or signals. The extracted biometric characteristics are represented as a set of points or vectors. The acquired modalities are stored as templates for future authentication, verification, and identification. Biometric features extracted from biometric samples are stored as a template, which is therefore called a biometric template [15].
II-C Blockchain for Biometrics
Blockchain and biometric systems can be integrated in various ways. Here, we discuss the process involved in a typical blockchain transaction intersected with the typical biometric workflow.
A transaction on the blockchain represents an interaction between the parties [16]. In the context of biometrics, recording the result of an authentication is considered a transaction. Common biometric operations, such as template creation, modification, deletion, and retrieval, are also considered transactions that are recorded on the blockchain[17]. The lifestyle of a biometric transaction on the blockchain is shown in Fig. 2.
This workflow begins when a user provides their biometric input, such as a fingerprint or facial scan, using a secure device. This raw biometric data is then processed to create a unique digital representation, known as a biometric template. To ensure privacy, the template is secured using the robust protection method. The protected template (PT) can be stored off-chain or on-chain. Given the sensitive nature and size of biometric data, it is typically stored in a secure off-chain database, while a cryptographic hash of the encrypted template is generated and stored on the blockchain. This hash serves as a reference to verify the integrity of the biometric data without exposing the actual template. A blockchain transaction is then created, including this cryptographic hash, and signed using the user’s private key to ensure authenticity. The signed transaction is transmitted to the blockchain network, where the nodes validate it by checking the digital signature and ensuring that it adheres to the protocol rules. Once validated, the transaction is included in a new block and added to the blockchain through the network consensus mechanism. For future verification, the user provides their biometric input again, a new template is generated and encrypted, and the system compares the hash of this new template to the one stored on the blockchain to confirm identity.
Integrating biometrics with blockchain technology encompasses various applications and can impact multiple modules within biometric systems. For example, Delgado-Mohatar et al.[18] explored implementing each component of a biometric template protection system on a blockchain. Biometric authentication is the most common application of biometrics. In this context, Lee and Jeong[19] proposed an end-to-end protocol consisting of enrollment and authentication called BDAS, based on blockchain, as shown in Fig. 3.
Figure 3: An overview of a biometric authentication system using blockchain. (Image taken from[19].)
The BDAS enrollment process involves (1) capturing user biometric data to create a template, (2) splitting this template into three fragments based on its timestamp, (3) identifying connected nodes, (4) selecting clients to store these fragments, each fragment stored in $\left\lfloor\frac{n}{3}\right\rfloor$ copies among $\left\lfloor 3\times\left(\frac{n}{3}\right)\right\rfloor$ clients, and (5) storing the assigned template fragments in the selected clients.
The authentication process comprises five steps: (1) a client captures biometric data to request authentication, (2) runs a LookUp smart contract to obtain the locations of the required template fragments, which (3) are returned; (4) the client retrieves these fragments through separate communications, (5) merges them into a complete template, and compares it with the stored data.
II-D Related Works
Several studies are related to the current survey:
In 2022, Ahmed et al. [20] provided an in-depth analysis of how blockchain technology is revolutionizing identity management systems (IdMs), particularly through the lens of Self-Sovereign Identity (SSI). Their study explored the SSI ecosystem, emphasizing the role of decentralized identifiers and verifiable credentials in empowering users with control over their personal data.
In the same year, Rathee and Singh [21] analyzed how blockchain technology can enhance IdMs by addressing security and privacy concerns. Their study showed that blockchain-based identity systems outperform traditional centralized systems in terms of transparency, security, and ease of use for both individuals and organizations. Similarly, Ngo et al. [22] provided a systematic overview of blockchain-based IdMs, classifying them according to their contributions, application domains, identity types, and research methodologies. Their study focused on papers from four databases (IEEE Xplore, ScienceDirect, ACM, and Springer Link) published between 2009 and 2022, and proposed general solutions for the application of blockchain to identity management.
The study that aligns the most closely with the objective of our survey is the work published by Sharma and Dwivedi [23]. They examined various applications of blockchain in biometric systems, including biometric template storage, identity management, and authentication processes. However, their analysis can be criticized for its superficial treatment of the subject matter, lacking depth in its exploration of the technical and legal complexities involved.
Although Rathee and Singh [21] mentioned Civic [24], which claims to be GDPR compliant, all these related works lack any assessment of legal or GDPR-based regulations regarding the combination of blockchain and biometrics.
III Combining Blockchain and Biometrics: Literature
Biometric systems can greatly benefit from the functionality and features offered by the blockchain. In this section, we argue for the strengths and weaknesses of key aspects that play a pivotal role in the integration of blockchain and biometrics.
III-A Scalability
III-A1 Storage
All blockchain-based applications face an inherent challenge with the size of blocks [25]. In existing scenarios, all transactions that are processed through the network have to be stored on the blockchain. In the blockchain network, all nodes are responsible for storing and maintaining the entire transaction record. With the current size and growing storage space of two prominent public blockchains, Bitcoin and Ethereum, exceeding 200GB [26], scalability has become the main barrier to adopting the blockchain in all applications, including biometrics. Multiple solutions have already been proposed, regardless of the application, to address this challenge. The first solution is to remove old recorded transactions and keep the balance of non-empty addresses using a database called the account tree [27]. This approach obviates the need for storing all transactions by keeping the hash of previous transactions to check the validity of new transactions.
The second solution would be to re-design the blockchain for the next generation of Bitcoin (Bitcoin-NG) [28]. The driving force behind this remedy is to divide conventional blocks into two separate parts: the key block and the microblock. In this revision, miners would compete for the key block to become the next leader who is eligible for generating a microblock wherein they store new transactions. Nevertheless, this solution suffers from some limits. For example, in terms of biometric verification, it needs to be lightweight due to the high utilization rate of this application. In case of biometric verification, if the history of all transaction are stored on-chain, it is essential to take care of storage obstacles. It is important to mention that it is considered risky by legal reasons to directly store raw biometric data or their extracted embeddings on a public blockchain. Thus, when we refer to storing biometric data, it usually means storing protected biometric templates. Lastly, the third solution, being considered by the most current literature [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] to overcome scalability issues with storage limitations [17], also applicable to biometrics, is to record blockchain transactions (biometric data) off-chain (any storage located outside of public blockchain). This solution which does not guarantee the immutability of the biometric data is divided into two approaches: (i) recording the hash of the biometric data (hash of the protected template) inside the public ledger and storing the biometric data itself off-chain in external storage systems, or (ii) recording only these hashes in a binary tree data structure called Merkle tree [26]. The key difference between these approaches lies in how they securely store content. The latter approach ensures data integrity through a recursive process of self-construction. The latest advancement in off-chain storage methods, designed to maintain the decentralization of public blockchains, is the use of IPFS [40]. The InterPlanetary File System (IPFS) is a distributed file system meant to be used for connecting computing devices with the same data structure. Like blockchain, it has no single point of failure where nodes do not need to trust each other. Yet, regarding the combination of blockchain and biometrics, most of the existing works suggest recording the hash of biometric data on-chain and storing the data itself on IPFS. Though this kind of storage introduces some deficiencies as the hash of data limits some biometric applications. For example, concerning biometric matching, extracted features are transformed into templates using different protection methods [41]. Therefore, storing the hash of templates inside a block makes it infeasible for biometric comparison function. Off-chain storage gives this opportunity to store the biometric data itself in addition to the hash of it. Hence, regarding biometric comparison, it can be done using a smart contract on off-chain networks. Currently, there are four recording approaches with regard to this combination in the existing literature: (i) Store biometric data as it is without any pre-processing on-chain (directly on the blockchain), (ii) Store biometric data off-chain on a side-chain (outside of the blockchain) and the hash of it on-chain, (iii) Store biometric data in distributed file systems (e.g., IPFS) and the hash of it on-chain, (iv) Store only the hash of biometric data on a Merkle tree. The pros and cons of each approach are given in Table I.
TABLE I: Current recording approaches for combining blockchain with biometrics found in literature [34, 17]: (i) Store biometric data as it is on-chain (on the public blockchain), (ii) Store biometric data off-chain on a side-chain (e.g., a private blockchain or any kind of side-chain) and the hash of it on-chain, (iii) Store biometric data in distributed file systems (e.g., IPFS) and the hash of it on-chain, (iv) Store only the hash of biometric data on a Merkle tree.
Approach Pros Cons
(i) Can be used for almost all biometric applications Taking advantage of all characteristics of a public blockchain Storage space in the current public blockchain is expensive Slow (not suitable for real-time use-cases) Wouldn’t comply with GDPR
(ii) Doesn’t have the scalability limitations of public blockchains (cheaper, faster) Since the data is stored off-chain, it’s not possible to take advantage of public blockchain characteristics like immutability, transparency and decentralization
(iii) Doesn’t have the scalability limitations of public blockchains (much cheaper, faster) Decentralized networks like a public blockchain It’s a distributed file system, not a blockchain Like storing data in public blockchain it is subjected to legal issues Need two storage locations (public blockchain and IPFS) The hash of data is not useful for all biometric applications
(iv) Doesn’t have the scalability limitations of public blockchains (less storage space, faster) Very limited in terms of the biometric application
III-A2 Latency
Writing new blocks on-chain is quite slow and completely infeasible for being adopted in real-time applications. For example, in terms of the most popular cryptocurrency, the required time to establish a new block in the Bitcoin network is 10 minutes on average, making it limited to processing up to 7 transactions per second. The same scalability limit lies ahead of the second generation of blockchain systems such as Ethereum with a limit of 15 transactions per second. As a result, biometric applications that require real-time responses, such as most verification and identity management systems, cannot integrate with these blockchain networks. Off-chain storage has also been proposed to solve the latency problem [42]. In this case, IPFS can address the content quicker, more reliably, and more easily by locating the stored data using the file information. Although off-chain storage helps reduce latency compared to on-chain solutions, it still imposes limitations on certain biometric applications (e.g., biometric authentication, surveillance, e-commerce), particularly similar to the storage challenges previously mentioned.
III-A3 Computation cost
It is discussed that another reason to suggest using off-chain storage is attributable to the total cost of the computation on-chain. The main biometric operations that impose arithmetic costs on the blockchain network involve setup, template transformation, encryption, decryption, upload (read, write), updating the smart contract, and calculating distance. However, the exact computational cost of biometrics in the blockchain is highly dependent on the template acquisition method [32], the underlying consensus protocol, and the operations executed by the contract [30]. For example, the initial cost for reading and writing 1KB of data on the Ethereum network is 6400 and 640000 gas units, respectively [17]. Therefore, storing 1KB of data from the vector of face characteristics for the VGG-Face template on the Ethereum network costs 352912 gas. Gas is a unit to pay for transactions on the Ethereum blockchain denoted in gwei, where 1 ETH = $1*10^{9}$ (1,000,000,000) gwei. It is a measure to compute how much a transaction would cost. Therefore, at the time of writing this article, the storage cost for 1KB data of the face feature vector on a full on-chain storage scheme is approximately \$0.69 (1 ETH = \$1974). Due to the computationally intensive process of biometric applications such as biometric verification, the widely suggested solution for massive scalability appears to be to resort to off-chain storage (approaches ii and iii in Table I), regardless of the limitations.
III-B Smart Contracts
Smart contracts can automatically execute verification processes when predefined biometric conditions are met, eliminating the need for intermediaries. In addition, the use of smart contracts on the blockchain enables users to grant or revoke permission for the usage of their biometric data. When implementing biometric applications on a blockchain using smart contracts, it is essential to consider two key issues: First, it is reported that gas estimation at the source code level is a difficulty for smart contract development. Since biometric applications are resource intensive, the underlying smart contract must be developed in a resource-friendly way. Second, special attention is needed for smart contract security and bugs, especially considering the immutability and non-reversibility of transactions. The record on the blockchain, the execution and the management of biometric templates will be handled by the smart contract. Therefore, any existing security problem or bug within the smart contract leads to a catastrophe. For example, more than 60 billion dollars were stolen from the DAO, decentralized autonomous organizations that are entities to operate through smart contracts, due to the recursive call bug [43]. As an alternative solution to these problems, an off-chain smart contract has been proposed [44]. Unlike on-chain blockchains, they are executed outside the blockchain by only interested participants. Although on-chain protocols are better for low-cost computation and non-sensitive transactions, an off-chain smart contract is suggested for operations involving high-cost computation or private information of participants. However, it is better to exploit the advantage of an on-chain smart contract jointly for its auditability and trustable characteristics regarding any disputes about the execution of an off-chain protocol.
III-C Consensus
In the blockchain context, a consensus is a mechanism to reach an agreement between participants who do not necessarily trust each other in a distributed network without a central authority. The consensus protocol is responsible for the correct execution of the smart contract on the blockchain. In combination with biometrics, the consensus algorithm affects the latency to execute operations. With the emergence of the next generation of blockchain networks [42], equipped with fast consensus algorithms that maintain the security benefits of their predecessors, the need for off-chain protocols, at least to overcome scalability limits, will likely be eliminated. To this end, the third generation of blockchain systems is emerging that targets scalability limitations. As an example of these new generations of consensus algorithms, Stellar consensus protocol [45], which is based on a new type of federated byzantine agreement [46], can process 1,000 transactions/second, Ripple can handle up to 1,500 transactions/second, and Hashgraph, which takes advantage of a new type of asynchronous, byzantine fault tolerance can achieve above 250,000 transactions/second. Recently, an identity management method called DID (Distributed Identifiers) [47] has been introduced which transfers computation overload off-chain, decreasing demands for the main blockchain. Due to its decentralized nature and ability to enable trusted interactions, DID serves as a viable alternative for storing biometric templates on-chain. We will discuss DID further in Section V-B.
III-D Security Enhancement
Owing to its unique features, blockchain can potentially help improve the security of distributed networks. In terms of the combination with biometrics, while it offers a breakthrough, multiple concerns need to be addressed. First and foremost, this integration can not only help introduce a public key infrastructure (PKI) platform [7] to biometrics, but also helps to overcome the problem of a single point of failure with which conventional PKIs are entangled. In conventional PKI systems, if the Certificate Authority (CA) is not accessible, it is not possible to verify the authenticity of generated certificates. Due to its availability characteristics, blockchain addresses this shortcoming. Second, considering biometric verification, it provides a distributed storage system in which multiple transformations of the same biometric template can be stored in different places [48, 49]. In existing scenarios, some tried to increase blockchain security by generating biometric cryptographic keys to secure the private key in blockchain wallets [50] or creating an encryption scheme to sign blockchain transactions [51], while others took advantage of the blockchain to increase biometric system security. Regarding the first, since the private key is stored in the user’s wallet, the security of the private key is dependent on the security of the wallet password. Biometrics can also be used as an access control to the wallet in a key binding approach [52]. In terms of the latter, which is the focus of this survey, most studies have been carried out to secure biometric verification using blockchain as a means to provide decentralized auditability, availability, and immutability. Some studies suggest blockchain for biometric applications where transparency is required, such as public data management, electronic voting, and taxation [39, 35]. Since the blockchain is immutable and the recorded information persists there forever, the fate of stored biometric templates would be shifted to the security of the underlying cryptographic method. Concerning auditability, referring to the 51% attack [53], it is possible to exert influence by most of the nodes that colluded for a single purpose and contradict the facts already recorded by creating a fork. Furthermore, due to the immutability of the blockchain, it is not possible to discard an expired biometric template stored earlier on a public ledger unless $51\%$ of the nodes colluded. Regarding private blockchains, since the owner of the blockchain controls the majority ($51\%$) of the miners, it is practically possible to remove an expired biometric template by resetting the last block to an earlier block and continuing to mine from there. However, this process is costly, since removing a biometric template from the blockchain also deletes all subsequently added templates. Thus, the owner of the private blockchain needs to rewrite the history after the discarded block. A recently proposed solution for Bitcoin introduces a miner-redactable and scalable approach that reportedly eliminates the need for trustees, heuristics, or hard forks. Kiraz et al. [54] claim to achieve instant erasure of all transaction types by enforcing transaction “templates”. These templates ensure that only the non-executable portions of unlocking scripts can be modified, preventing unintended alterations. In particular, a significant portion of such nonexecutable data consists of nonfinancial information such as biometric templates or references to them within the Bitcoin network, which are stored as OP_RETURN data. In 2017, reports indicate that 86.8% of Bitcoin’s non-financial data falls into this category. Furthermore, this approach eliminates the need for modifications in child transactions after redaction, effectively resolving the issue of “cascade of changes” found in $\mu$Chain [55], a redactable blockchain designed for Ethereum.
III-E Privacy Preservation
In general, each node in the network must execute and validate a smart contract on the chain. To this end, all relevant transactions are visible to the entire blockchain network, resulting in privacy leakage. Privacy leakage has been shown to occur on the blockchain even when users rely only on their public key to create a transaction [56]. Since the blockchain is transparent, it is possible to link transactions made using the same public key to an identity. When it comes to the processing of biometric data for very sensitive applications such as e-health, it is essential to address data privacy and anonymity issues [57]. In addition, since public keys are bound to biometric templates, linking a public key to a real identity may result in permanent leakage of the biometric data. Although homomorphic encryption can be used to keep and process biometrics in cipher for fixed length templates [58, 41], and other privacy-preserving methods can be applied for variable length templates [59], there is no guarantee that the protected template will remain unbroken in the future.
Bitcoin provides only weak identity privacy. To overcome this limitation, several privacy coins have been introduced, including Dash, Zerocoin, Monero, Zcash, Zether, Grin, Verge, and MimbleWimble, among others [60]. These cryptocurrencies aim to enhance transaction anonymity and protect user identities, making them relevant to privacy-focused, permissionless blockchains designed for identity privacy.
One of Bitcoin’s main privacy vulnerabilities is that data miners can link individual transactions and track user activities. A common solution to this issue is the use of Bitcoin laundering services, also known as mixers, which mix multiple users’ bitcoins to obscure transaction histories. Hence, one of the most common recommendations is to employ a laundry service such as CoinShuffle++, which facilitates coin mixing to enhance privacy [61].
Zerocoin (formerly Zcoin, XZC) is a proposal to extend Bitcoin with anonymous transactions. This extension would have functioned as a money laundering pool, where bitcoins were temporarily pooled in exchange for a temporary currency called zerocoins. Although laundering pools are already used by various currency laundering services, Zerocoin aimed to integrate this functionality at the protocol level, eliminating the need for trusted third parties. Ensured anonymity using cryptographic principles while still recording transactions within the Bitcoin blockchain [62].
Monero (XMR) was the first blockchain designed with the privacy and resistance to censorship as the core principles. Every Monero transaction is anonymous by default, unlike Bitcoin, where transparency is the norm. Monero uses the CryptoNote protocol, which employs ring signatures and one-time keys to obscure the origins and destinations of the transactions. This approach ensures strong privacy, making Monero one of the most widely adopted privacy coins with a high market capitalization [63].
Zcash (ZEC) is another major privacy coin that offers confidential transactions. Unlike Monero, where privacy is enforced by default, Zcash provides users with the option to choose between shielded (private) and transparent (public) transactions, granting greater control over privacy preferences [64].
These privacy coins demonstrate how permissionless blockchains can be leveraged for identity privacy. By integrating cryptographic techniques such as ring signatures, zero-knowledge proofs, and coin-mixing mechanisms, these blockchain solutions provide individuals with enhanced control over their biometric data, mitigating the risks associated with traceable transactions.
More sophisticated general privacy-preservation mechanisms such as differential privacy, to our knowledge, have not yet been adapted and studied to blockchain.
Finally, biometric-specific privacy preservation methods (e.g., obscuring sensitive elements on biometric embeddings [65, 66]) can be applied directly.
III-F Literature Categorization
Table II provides a comparative overview of existing works that integrate blockchain technology with biometric systems, focusing on key technical aspects. The comparison highlights crucial parameters such as the biometric type (modality), the blockchain type (permissioned vs. permissionless), framework, storage approaches (on-chain vs. off-chain), integration purpose, blockchain network, and key technologies.
This comparison aims to identify the strengths and limitations of different approaches, providing insight into how blockchain enhances biometrics for the given integration purpose. By analyzing these technical attributes, researchers and practitioners can better understand the trade-offs involved in various implementations and identify potential areas for future improvement.
TABLE II: Comparison of existing works in terms of key technical information (N/S stands for Not Specified).
Scheme Biometrics storage location Biometric modality Framework Blockchain Type Integration purpose Blockchain Network Key technologies
[67] On-chain Fingerprint Biometric authentication Permissionless Safe storage of biometric vectors in blockchain Ethereum MetaMask, Cross-numbering
[68] On-chain Fingerprint Biometric authentication N/S Availability, immutability, decentralization, anonymity N/S N/S
[69] On-chain Fingerprint Biometric authentication N/S Providing availability Ethereum N/S
[70] On-chain Fingerprint Biometric authentication Permissionless Privacy preservation Monero Public key cryptography, Zero knowledge proofs
[71] On-chain Fingerprint Electronic voting system Permissionless Cost efficiency, trust Bitcoin Elliptic Curve Digital Signature (ECDSA)
[72] On-chain Fingerprint, face Secure biometric recognition Permissioned Combat tampering, provide fault tolerant N/S Multi-PIE, MEDS-II, Public key cryptography
[73] On-chain Fingerprint, face Biometric identification Permissionless Providing privacy Bitcoin Elliptic curve, RSA
[74] On-chain Fingerprint, iris Identity management Permissionless Long-term security, post quantum security Bitcoin directed acyclic graphs (DAG)
[75] On-chain Fingerprint, finger vein Multimodal biometric N/S Scalability, security N/S N/S
[76] On-chain Finger vein Biometric authentication N/S Data integrity and availability N/S Steganography, Gaussian smoothing, segmentation
[77] On-chain Fingernail Fingernail analysis management system N/S Privacy, immutability and trust N/S HOG, LBP
[78] On-chain Face Biometric authentication N/S Secure communication N/S Shift-AES
[79] On-chain Face Biometric authentication N/S Availability, immutability N/S Fuzzy extractor and secure sketch
[80] On-chain Face Biometric authentication Consortium blockchain Cross-domain authentication N/S N/S
[81] On-chain Face Reliable food logistics N/S Security, scalability ad privacy N/S N/S
[82] On-chain Face Secure vehicle data-sharing N/S Further security N/S N/S
[17] On-chain Face, signature Biometric templates storage N/S Availability of Biometric templates Ethereum Ropsten Ethereum tesnet, Biosecure DS2
[83] On-chain Signature Biometric signature N/S N/S N/S N/S
[84] On-chain All modalities Authentication Permissioned Providing traceability and reducing fraud Hyperledger fabric N/S
[19] On-chain N/S Biometric authentication Permissioned Decentralization, auditablity, biometric template security Ethereum Geth/Solidity
[85] On-chain N/S IoT safe access N/S Transparency N/S N/S
[86] On-chain N/S Biometric based signature N/S Providing PKI mechanism in blockchain N/S Fuzzy signature
[32] Off-chain (IPFS) and on-chain Fingerprint Identity management Permissionless Providing Immutability to biometrics Ethereum IPFS
[87] On-chain and off-chain Fingerprint, face Biometric authentication Permissionless Transparency, convenience N/S Decision tree
[29] Off-chain: on-device Fingerprint Biometric authentication Permissionless N/S N/S ORB descriptor
[33, 34] Off-chain Fingerprint Blockchain-based Identity management Permissionless Prevent credential transfer Bitcoin Fuzzy extractor, fuzzy vault
[35] On-chain Fingerprint, iris Electronic document identification N/S Traceability, security, scalability N/S N/S
[36] On-chain Finger vein Library management N/S Availability N/S N/S
[31] Off-chain: on-device Face Identity management N/S Cost effective identity binding N/S TrustZone
[88] Off-chain:IBM’s Blockchain Face Identity management Permissioned Security and privacy Hyperledger Fabric Bloom Filter
[37, 38] Off-chain: wallet Iris Blockchain authorization N/S Addressing double spending N/S IIF algorithm, Spatial histogram
[39] Off-chain and side-chain Iris Online voting system Permissionless Decentralize and tamper proof storage of biometric templates Ethereum Meta mask, Gabor filter, Rutovitz crossnumbering
[30] Off-chain: IPFS N/S Biometric authentication N/S Cost effective identity binding N/S IPFS, homomorphic encryption
[89] N/S Face Authentication Permissioned N/S N/S Fuzzy vault
[90] N/S Face Authentication N/S Privacy, Revocability N/S random distance method (RDM)
[51] N/S Face Transaction authorization N/S Addressing double spending problem in blockchain N/S HOG face detector, BNIF algorithm
[52] N/S Face Key management N/S Securing wallet keys N/S HOG, Fuzzy vault
[50] N/S Behavioral biometric Key management N/S Continuous authentication N/S 12-layer CNN network based on keras
[91] N/S Psychological traits Multimodal biometric authentication Permissioned Liveliness detection before authentication N/S SaaS
[26] N/S N/S Biometric template protection N/S Immutability, accountability, availability N/S N/S
IV Blockchain and Biometrics: GDPR Elements
As mentioned above, in contrast to traditional centralized database systems, blockchains allow "selective data exchange" with certain stakeholders, which can reduce the chance of identity theft and privacy problems [92, 93]. However, it brings about many other problems regarding the applicability of the GDPR. The analysis below is categorized according to five relevant parameters: (i) biometric data storage, (ii) accountability, (iii) immutability, (iv) anonymization, and (v) data protection by design and default. These are further addressed in the following, taking into account the technical analysis.
IV-A Biometric Data Storage
This parameter is related to the questions as to how and where the template is stored. Interdisciplinary research [94] in 2009 concluded that biometrics could be invasive to privacy or improve to privacy depending on the design and safeguards. In addition to the research analyzed above, blockchain surveys show that depending on the types of network and storage, these systems might have different levels of GDPR compliance [95]. These two storage criteria are often suggested as key determinants of GDPR compliance level, whereby a helpful distinction can be made based on on-chain and off-chain storage.
IV-A1 Off-chain storage
Central storage solutions are not considered privacy-friendly, as observed by the European Data Protection Supervisor [96]. Off-chain storage can be decentralized and such strict decentralization methods are considered more privacy friendly [97] as the data controller cannot access the biometric templates and only receives the comparison result. However, data security is a crucial part of the data protection law, and this type of decentralization may also be susceptible to attacks. As explained in various reports, such as by the BSI, local storage of biometric data enabled by software, for example, rich iOS of a mobile handset, cannot be regarded secure [98]. ENISA further observed that multiple biometric templates can be stored in Touch or FaceID and can belong to more than one person [99].
A common misunderstanding in the technical papers analyzed is that off-chain biometric data storage is, by default, GDPR friendly. However, as explained, off-chain storage itself does not render the processing compliant or secure. This implies that even if the biometric data are stored off-chain, personal data linked to it, such as a user ID, public keys, and the content of the transactions, are usually still stored on-chain. In such cases, even if the public keys, e.g., hashed identifiers, change in every transaction, they still qualify as personal data under the GDPR, as the natural persons remain identifiable.
IV-A2 On-chain storage
On-chain processing, that is, the processing "located, performed, run inside a blockchain system" [100] of the special categories as such, while risky for the reasons we will discuss in detail later, is not per se against the GDPR, as long as the processing is lawful. For instance, in cases where the explicit consent of the data subject or another legal and legitimate ground under Article 9 is relied on, and the other GDPR requirements are met, this processing by itself would not infringe the GDPR. However, on-chain processing of personal data is not recommended for the reasons explained in further detail.
IV-B Accountability
As explained previously, one of the most critical problems identified in the legal literature is how to allocate responsibility for data processing activities in the face of the polycentric nature of blockchains (i.e., accountability) [101, 102, 1]. Although private blockchains are less inclined to have such problems, the real benefits of blockchains, namely transparency, can be seen in public blockchains. That is why it is not surprising that identity management systems based on blockchain (and biometrics) are positioned to run their projects on public blockchains. This can be seen as a general conceptual issue with respect to public blockchains and requires more legal guidance and research. Hence, unless more legal clarity is established on the issue, processing via public blockchains is risky, as the entity responsible for GDPR compliance is unclear.
This general concern about blockchains is amplified when it comes to biometric data processing, as they are universally distinctive enough to identify people. Furthermore, biometric data are not chosen by individuals and cannot be reissued. Thus, biometric data are considered more sensitive and their combination with blockchain technology can result in long-term linkages between personal data and biometric IDs [103].
Riva [104] suggests that an accountable agency should be assigned top to bottom to eliminate the accountability issue. The decentralized nature and benefits of blockchains are severely undermined in this model. The French Data Protection Authority (CNIL) asserts that professional or commercial activity should be the main parameter when assessing controllership [105]. Although this may be a practical solution, the jurisprudence of the European Court of Justice holds that the controller should be determined through a factual analysis, which means that a factual influence is necessary for an entity to qualify as a data controller [106].
IV-C Immutability
Immutability as the main attribute of blockchain may compromise the privacy of data subjects, particularly where their biometric data are processed [107]. Article 17 GDPR gives effect to the demands of individuals (data subjects) to have their data deleted without undue delay. As we just stated, biometric traits are unchangeable and biometric templates are, in theory, irrevocable. Therefore, any blockchain design that envisions raw or hashed on-chain storage of biometric templates would exacerbate the data protection risks. However, the same article also defines the exceptions to the right that are important to keep in mind when experimenting with blockchains and personal data. These exceptions demonstrate that the right to be forgotten is not an absolute right, nor is the right to privacy and the right to the protection of personal data. These exceptions are as follows.
  • Exercising the right of freedom of expression and information. The right to freedom of expression and information is one of the rights that might conflict with the right to protection of personal data.
  • Compliance with a legal obligation which requires processing by the Union or Member States law to which the controller is subject for the performance of a task carried out in public interest. For example, research [71, 39] proposing an electronic voting system can be considered in the scope of this exemption.
  • For the performance of a task in the exercise of official authority vested in the controller. This exception can be the case if the controller processes personal data on-chain to carry out a legal obligation, such as demonstrating compliance with the tax regulation.
  • Reasons of public interest in the area of public health (e.g., processing personal data on-chain to facilitate compliance with the COVID-19 restrictions at the borders or public activities).
  • Archiving purposes in the public interest, scientific or historical research purposes, or statistical purposes. This exception provides a broad freedom to scientific projects processing personal data on-chain. However, other exceptions must be taken into account when products are placed on the market. For blockchain-based personal data processing, this exemption can be problematic, as it might be very difficult to identify a certain phase when scientific research ends and the product is placed on the market.
IV-D Anonymity
Under GDPR, encrypted personal data are considered pseudonymous in principle, not anonymous. Pseudonymity refers to the cases where personal data are linked to unique identifiers, the processing of which is still subject to the GDPR. Article 29 Working Party states that ‘removing directly identifying elements in itself is not enough to ensure that identification of the data subject is no longer possible. It will often be necessary to take additional measures to prevent identification, again depending on the context and purposes of the processing for which the anonymized data are intended’ [108, 109]. However, most blockchain research analyzed uses the term anonymous, for the data that are hashed or encrypted. This approach might create an incorrect understanding and implementation of the GDPR, where the responsible entities consider that the processing operation is not within the scope of the GDPR. It is specifically risky when it comes to hashed biometric templates, since these data are considered personal data [110].
IV-E Data Protection by Design and by Default
For all the four problems explained above, another very relevant principle is data protection by design and default provided by Article 25 (1) GDPR. This principle is highly inclusive and starts to apply when the data controller determines the means of processing. Arguably, the developers of a blockchain-based biometric system can also be responsible. However, there are many arguments against this idea [111, 112]. However, the principle requires the data controller to take into account the state of the art, the cost of implementation, the nature, scope, purposes of personal data processing, and the risks and severity for the rights and freedoms of the data subjects. Biometric data processing is by nature considered risky. However, as mentioned, blockchain is still emerging and thus is considered a new technology whose impact on fundamental rights is unknown. Therefore, the processing of biometric data via blockchain is considered an innovative use of a large-scale operation that involves sensitive data where several datasets are matched or combined, and shall be subject to a Data Protection Impact Assessment (DPIA), under Article 35 GDPR[109].
A common misunderstanding in this regard is that scientific papers use the term privacy, which refers solely to the security of the data. Privacy also encompasses informational privacy, that is, data protection. As explained, data protection requires an array of technical and organizational measures, such as the principle of security that requires personal data to be kept confidential. This further means that personal data should be protected against accidental, unauthorized, or unlawful access, use, modification, disclosure, loss, destruction, or damage, according to Recital 39 of the GDPR. For example, using encryption techniques to pseudonymize personal data is considered an important element in data protection by design and default; however, other provisions of the GDPR should be respected when carrying out such processing, as pseudonymization itself does not render the processing GDPR-compliant.
V Future Directions and Use Cases
V-A Blockchain and Biometrics: Future Directions
According to the results of previous research discussed in previous sections, we provide some suggestions and recommendations for future directions toward the combination of blockchain and biometrics that should be taken into account when designing a technolegal framework that incorporates blockchain in biometrics.
V-A1 Biometric data should not be stored on a public ledger
As discussed in section IV, before having clarified the legal uncertainty on accountability, it is currently advised not to store biometric data on public blockchains. A public blockchain is an immutable storage that is accessible to anyone [100]. This means that once the data are published on the blockchain, it remains there forever. Considering the fact that biometric data contain information about our permanent identity and they are irreplaceable in case they leak. There are no guarantees that encryption mechanisms or biometric template protection methods remain completely safe in the future. Biometric data are discouraged in any form on a public blockchain.
In addition, from an efficiency point of view, blockchain storage is expensive both computationally and economically. In particular, reference [17] concluded practically that storing biometric templates on-chain or as hashed data is not cost-efficient. In addition to that, references [30] and [32] resorted to IPFS to solve the efficiency problem. Although using a distributed filesystem offers availability, it does not provide the security benefits that a blockchain provides, such as trust. Approaches (ii) and (iii) in Table I are proposed to address this concern.
V-A2 The integration should not be used for real-time scenarios
Currently, the process of publishing a new block and its confirmation time on a public blockchain is very slow. The latency of existing blockchain networks rules out real-time applications like biometric authentication and online payment. The most popular and widespread blockchain network, Bitcoin, is limited to publishing 3-4 transactions per second. At the time of writing this article, the time required for a Bitcoin blockchain network to issue a new block is 10 minutes. And this is without considering the confirmation time, which is variable and depends on the status of network. The second most popular public ledger, Ethereum, which is an open source and practical framework of most research implementations, is limited to 15 transactions per second with an average block publication time of 15 seconds [42].
V-A3 The integration should not be used for storing substantial data
At present, blockchain storage is expensive and is not cost-effective for being used in applications that work with large volumes of data. Since all full nodes keep a copy of all transactions, keeping large amounts of data on-chain results in increased storage costs and processing expenses. Therefore, the scalability limit of the blockchain should be taken into account before using it as permanent data storage.
V-A4 The integration should have economic justification
The life cycle of permissionless blockchains depends on the service of the miners who maintain the system. It should be economically interesting for these nodes to carry out their maintenance. To this end, the network pays a reward as a fee to miners who allocate their computing resources to maintain the network. For example, in Ethereum, this fee is called gas, and the required amount of gas to run the instructions inside the network differs according to their computational complexity [26]. Therefore, running smart contracts on future frameworks that combine blockchain and biometrics should be economically justifiable.
V-A5 The integration should not rely on blockchain for privacy
One of the intrinsic characteristics of blockchain is providing pseudonymity. This feature enables entities to exchange data with each other without knowing the identity of each other. As a result, some existing integration [68, 77] relies on this attribute to offer protection of privacy in biometric systems. However, blockchain entities are not fully anonymous, as they need to use a public key to send and receive transactions. In fact, such secrecy can be provided by encryption without the need to use a blockchain. In contrast, since the permissionless blockchain is transparent and the history of all transactions is accessible to anyone, users can be deanonymized using profiling [113]. Therefore, any framework that incorporates blockchain into biometrics should not consider blockchain as a privacy-preserving add-on.
V-B Combining Blockchain and Biometrics: Use Cases
So far, the most attractive applications for integrating blockchain and biometrics have been biometric verification and identification. Although these are conventional and widespread biometric applications, they are not necessarily the most fruitful use cases. After delving into existing studies, we conclude that using blockchain was not strictly necessary for many of these applications and was just added because of the recent blockchain hype. For example, it is mentioned in reference [81] that it is not clear to them why the blockchain is better than any regular storage system, yet they still use it. In the following, we propose some highly potential use cases for integrating blockchain and biometrics considering the pros and cons discussed in the rest of the paper.
V-B1 Using blockchain as computing power for biometrics
The first consensus protocol employed in the blockchain network is Proof-of-Work (PoW), used in the Bitcoin network [114]. PoW is a complicated and resource-intensive strategy to select the miner who wins the competition, being entitled to publish the next block on the network. The difficulty level of this competition is continuously adjusted by the network to maintain trust and preserve the immutability of the blockchain. Despite these advantages, PoW is not economically efficient or environmentally friendly as it consumes substantial electrical power. When a block is published, only one node (miner) is rewarded for allocating its computational power, while the resources of other nodes who lose the competition are wasted. To remedy this issue, Proof-of-Stake (PoS) has been proposed [115]. However, PoS also has its problems regarding security threats and stakes issues [43]. A fruitful use case that benefits both blockchain and biometrics concerning PoW might be to recruit or rent the computational power of miners. Biometric recognition can be a computationally expensive task, especially in relation to the model training phase. It is highly profitable to harness the computational power of miners in PoW-based blockchains to carry the burden of biometric computation for tasks that existed in different stages of a biometric system, such as training a biometric model, generating a protected template, and comparing biometric templates. In particular, it seems that the blockchain could be a profitable platform for both clients and model owners to operate a federated learning system [116] for biometrics. Therefore, in the proposed consensus algorithm, those nodes who dedicate their computational power to the benefits of the blockchain service for the biometric system are rewarded according to their contribution.
V-B2 Using blockchain as a platform for distributed identity management
With the rapid growth pace of the digital world, it is very important to ensure appropriate access to resources available by our different identifiers (e.g., phone number, email address, domain names) across heterogeneous technology environments and platforms. Recently, an architecture called Decentralized Identifiers (DIDs) [^1] https://www.w3.org/TR/did-1.0/ has been introduced as a new type of identity management system that allows for verifiable decentralized digital identity. According to the authors’ definition, DIDs are URIs that associate a DID subject with a DID document, allowing trustable interactions associated with that subject. They have been designed to be decoupled from centralized registries, certificate providers, and identity suppliers. Figure 4 shows an example of DID.
Figure 4: A simple example of a decentralized identifier (DID) and DID document.
This structure benefits the biometric system by fully controlling the subject’s identifiers with an individual biometric identifier and independently of different heterogeneous providers. Thus, biometrics could significantly improve the currently distributed digital identity schemes based on blockchain [26]. In terms of on-chain challenges, DIDs help to transfer most of the processing load from the main blockchain to the off-chain network. A DID document consists of three fundamental parts: an authentication key, an authentication method, and a service endpoint. Whoever has access to the key is entitled to access the DID document. An example of this use case is the work of [90], which proposed a blockchain-based biometric authentication system where biometric data is decentralized and managed using DID and DID documents. We recommend that future research use biometric data only in combination with similar specifications, such as off-chain storage, to minimize both the quantity and sensitivity of personal data stored on-chain.
V-B3 Using blockchain as a foundation for biometrics-based PKI mechanisms
Public Key Infrastructure (PKI) comprises the policies, processes, server platforms, software, and workstations necessary to administer public-private key pairs and digital certificates. Its primary functions include issuing, maintaining, and revoking certificates that associate a public key with a specific entity, enabling secure communication and authentication within network environments[117].
Integrating biometrics with PKI is an emerging approach aimed at enhancing security and user authentication. However, the adoption of biometric-based PKI systems is not yet widespread. Traditional PKI relies on passwords or hardware tokens for authentication and is managed by centralized authorities, making it vulnerable to single points of failure and inefficiencies. Incorporating biometrics can improve security by linking cryptographic keys to unique physiological traits. In contrast, blockchain-based PKI decentralizes key management in a network of nodes, reducing dependence on a central authority and improving resilience[118]. Despite this potential, several factors contribute to its limited adoption.
  • Cost and complexity: Implementing biometric authentication within PKI systems requires specialized hardware and software to capture and process biometric data, leading to increased costs and complexity compared to traditional methods[117].
  • Lack of trust: PKI has limitations in establishing trust due to concerns about relying solely on keys to represent identities [119].
  • Certificate’s life-cycle management: The handling of expired certificates is a major problem that exists in all PKI mechanisms.
  • Standardization: The lack of universally accepted standards for integrating biometrics with PKI poses challenges to interoperability across different systems and platforms.
Blockchain’s decentralized nature reduces the need for centralized authorities, potentially lowering infrastructure costs and simplifying system management. It also ensures that all participants, regardless of their location or system, follow the same protocols that facilitate the adoption of common standards across different platforms and systems.
World Network (formerly Worldcoin) and Humanity Protocol are examples of biometric-based PKI systems that use blockchain technology. World Network employs the biometric iris scan depicted in Fig. 5 to verify human identity, creating private crypto tokens [120, 121] for applications such as universal basic income and fraud prevention in virtual communications[122]. (Note that iris biometrics can also be captured with standard smartphone cameras [123], and iris biometrics, like other biometrics, are also susceptible to potential attacks and vulnerabilities [124].) Similarly, the Humanity Protocol uses palm scans [125] to authenticate online accounts, addressing issues such as bots [126, 127, 128], fake accounts, and online fraud, using blockchain and biometrics for secure identity verification[129].
Figure 5: Iris-scanning Orb of the identity-verification project Worldcoin, now called World [122].
V-B4 Using blockchain as a multimodal biometric authentication system
Blockchain helps the multimodal biometric system by providing a decentralized, secure, and privacy-preserving mechanism to manage access to the distributed stored encrypted biometric templates. Addresses the vulnerabilities of centralized systems by eliminating the single point of failure, reducing the risk of data breaches, improving system availability, and empowering users with control over their biometric data through smart contracts. The results of each modality can be published and linked on the blockchain. As a result, several applications can authenticate the same individual simultaneously. In addition, most of the time, the underlying biometric hardware that captures user biometrics is not made or maintained by the same party. Therefore, publishing the authentication result for each modality on the blockchain helps to eliminate inconsistencies between the underlying infrastructure of each hardware. These results can be linked using an identifier, for example, using DIDs. Therefore, it is possible to link the biometric modalities of individuals using DIDs on the blockchain. Although standardization and risk mitigation efforts, such as DIDs, are appreciated, it should be kept in mind that DID documents can still be linked to individuals, and hence they are considered personal data under legal rules like the GDPR. Another practice is the distributed storage with IPFS. To this end, a recent study [130] stored decentralized fuzzy vaults of facial biometric images and dorsal hand images on IPFS and their address on the Ethereum Virtual Machine (EVM)-based blockchain network. Once the addresses of the fuzzy vaults are recorded on the blockchain, they become immutable, making it extremely difficult for malicious attackers to manipulate or delete this crucial information without consensus from the network participants.
V-C Suggestions to Mitigate Legal Risks
First of all, we suggest that biometric data and any other personal data, such as name, nationality, etc., should not be stored on-chain until further research confirms that protective measures are sufficient. [131, 105].
Furthermore, to mitigate data protection risks, the objective of processing operations must be clearly stated and compared to the applicable provisions outlined above to comply with GDPR. For instance, in cases where the purpose of the processing is in the public interest, e.g., electronic voting system, public health, or processing personal data on-chain to facilitate compliance with the COVID-19 restrictions at the borders, the exemptions to the right to erasure (Article 17 GDPR) could be relied on.
When it comes to research purposes, the GDPR provides broad freedom to scientific projects processing personal data on-chain. However, this exemption cannot be relied upon when products are placed on the market. Hence, we suggest that the objective of processing operations must always be clearly stated and compared to the applicable provisions outlined above to comply with GDPR. This is not a simple procedure because blockchains operate in numerous industries and stakeholders. Furthermore, blockchains can serve various functions, from smart cities to medical records storage, meaning that many stakeholders should take the necessary technical and organizational measures to comply with GDPR. This multistakeholder framework not only challenges the right to be forgotten, but also points to a meeting point for many purposes of different sectors. However, when we look at the reasons for incorporation in several studies analyzed [87, 68, 77, 32, 39] we see that the immutability of the blockchain is the purpose of integration. In our view, processing on blockchains requires an answer to erasure requests of data subjects. The right to erasure is not likely to be ensured, as the data registered in the blocks are very difficult and costly to delete, even if a court decision orders such, e.g., in the case of unlawfully processed data.
Finally, with respect to public blockchains, there is uncertainty as to which entity is accountable, that is, the data controller. In these regards, private blockchains seem to pose fewer issues, but do not provide the real benefits of decentralization.
V-D Recommendations for Interdisciplinary Research
To improve the understanding and implementation of data protection by design, it is crucial, first of all, to establish a standard and harmonized understanding of the key terminology between legal researchers and computer scientists. This recommendation explicitly regards the term anonymity as explained, which is often misunderstood by the scientific community. In addition, privacy-preserving technologies should not only be understood as a means of improving data security. Although highly important, data security is only one aspect of the data protection framework. Thus, when developing privacy-preserving technologies, technology should incorporate other data protection principles enshrined under the GDPR, such as purpose limitation, lawfulness, data minimization, storage limitation, and accountability. Additionally, data protection analysis should not only focus on generic blockchain architectures, but also on additional technologies, as each project utilizes different complementary technologies to increase the efficiency and security of their systems. We call for specific guidance to mitigate these risks, taking into account the sensitive nature of biometric data.
VI Conclusions
In this survey, we investigated the advantages and shortcomings of the mutual integration of blockchain and biometric recognition systems. The goal of this survey is to clear the path for this research line to further develop and to clarify some of the legal requirements while considering the technical status and advancements of this technology. To sum up, this combination can be favorable for biometrics as it provides a platform for PKI, distributed identity management, federated learning, and multimodal biometrics. On the other hand, we found that it contains risks for real-time applications or when used as a storage space for sensitive information such as biometric data. Nevertheless, with further advances in blockchain technology, some of these conditions may change. From a legal point of view, one shall take into account the sensitive nature of biometric data and the specific emerging risks out of this combination, such as accountability. It is important when experimenting with biometric data and blockchain, researchers keep both raw biometric data (images) and templates off-chain. Moreover, the purpose of the processing should be clarified for each granular processing operation, and the risks should be assessed in order to apply additional safeguards for a future-proof data protection design. Better precision of the purpose of the processing can also help understand whether the exemptions to the right to be forgotten (Article 17 GDPR) apply to that processing. Finally, considering the current literature and the insights of the authors, we provided some recommendations and future directions for combining blockchain and biometrics, such as achieving a harmonized language between legal and technical research.
Acknowledgment
This work has been supported by projects: PRIMA (ITN-2019-860315), TRESPASS-ETN (ITN-2019-860813), INTER-ACTION (PID2021-126521OB-I00 MICINN/FEDER) and BBforTAI (PID2021-127641OB-I00 MICINN/FEDER).